OpenStudy (anonymous):

5(2x - 7) = 33 - 3(6x + 4)

OpenStudy (anonymous):

10x-35=33-18x-12 56=28x x=2

OpenStudy (anonymous):

x=2 I used a calculator

OpenStudy (anonymous):

\[10x-35=33-18x-12\] \[10x-35=-18x+21\] \[28x=56\] \[x=2\]

OpenStudy (anonymous):

Thank you